# Forward-mode automatic differentiation for the tensor renormalization group and its relation to the impurity method

### v1

@@ Abstract
We propose a forward-mode automatic differentiation (AD) framework for tensor renormalization group (TRG) methods. In this approach, evaluating the derivatives of the partition function up to order $k$ increases the matrix-multiplication cost by a factor of $(k+1)(k+2)/2$ compared to computing the free energy alone, while the memory footprint is only $k$ times that of the original calculation. In the limit where the derivatives of the SVD are neglected, we establish a theoretical correspondence between our forward-mode AD and conventional impurity methods. Numerically, we find that the proposed AD algorithm can calculate internal energy and specific heat significantly higher accuracy than the impurity method at comparable computational cost. We also provide a practical procedure to extract critical exponents from derivatives of the renormalized tensor in TRG calculations in both two and three dimensions.
@@ II.3 Differentiation of the HOTRG
The derivatives of E,FE,\,F can be obtained through a differentiation of the SVD formula, which we leave for appendix B. In the practiacl calculation, we adopt a Lorentzian broadening $1/x\rightarrow x/(x^{2}+\eta)$ for derivatives of isometries following Ref. [59] to avoid divergence when there are degenerate singular values. In the Ising model which we use in this work, $\eta\rightarrow\infty$ limit corresponds to the vanishing limit of the derivatives of squeezers [^1] This is because initial tensor of Ising model has reflection symmetry, which yields squeezers to be isometry itself. In general squeezers, $\eta\rightarrow\infty$ doesn’t mean $\dot{E}=\dot{F}=0$ without taking derivatives of singular values to be zero.. The above procedure completes the formulation of the first-order response within the forward-mode scheme. The extension to second-order derivatives is completely analogous. By differentiating Eq. (15) once more with respect to $\beta$, we obtain the evolution equation for the second derivative $\ddot{T}^{(n+1)}$ as follows. $T^{(n)}$, $\dot{T}^{(n)}$, and $\ddot{T}^{(n)}$.
@@
Higher-order derivatives follow straightforwardly by iterating the same procedure. The update rule for $\dot{T}$ in Eq. (15) naively consists of four HOTRG-like contractions. However, we do not need to evaluate these four contractions separately. Instead, we use the fact that each contraction can be decomposed into a sequence of matrix multiplications along a common contraction tree. Rather than evaluating the four contractions independently, we differentiate the intermediate tensors that appear along this tree. With this procedure, the total matrix multiplication cost of computing derivatives up to first order is exactly three times that of the original HOTRG.
@@
The blue colored tensors indicate its second derivative. In this contraction, the bottleneck takes just three times as the HOTRG. In practice, it is not necessary to execute all contractions shown in the figures independently. Instead, by performing the bulk, first-order, and second-order contractions sequentially for each intermediate tensor, one can reuse the intermediate tensors that are common to Eqs. (II.1), (II.3), and (II.3), rather than recomputing them multiple times. As a result, the total cost of evaluating derivatives up to first and second order is exactly three and six times that of the original HOTRG, respectively. This scaling is unchanged for general tensor network contractions. See Appendix C.1 for details. We also note that the memory footprint is only $k$ times that of the original HOTRG, since we only propagate $T^{(i)}$ and its derivatives at every coarse-graining step; importantly, it does not depend on the graph depth $\log V$. Note that, to obtain the derivatives of the squeezers (isometries), we differentiate the SVD and apply the resulting rules to all contractions appearing in the original algorithm. This approach is applicable to a wide range of programming languages without relying on external AD packages, as it only requires implementing a simple update rule.
@@ III Numerical results
Next, we discuss the computational cost. We first measure the elapsed time of the bottleneck contraction part in Eq. (II.1) and its derivatives, or impurity update rule. Figure 2 shows the elapsed time of the bottleneck contraction in HOTRG for the forward-mode AD computations up to first- and second-order, and for the corresponding first- and second-order impurity tensor computations. These correspond to Eqs. (II.1), (II.3), and (II.3) for the forward-mode AD, and to Eqs. (II.1), (30), and (31) for the impurity method. For the forward-mode AD, we employ contraction tree based differentiation described in Appendix C.1. All benchmarks are performed on a single CPU using randomly generated dense tensors, and the bond dimension is varied from $D=10$ to $70$. The vertical axis shows the elapsed time normalized by that of the original HOTRG contraction in Eq. (II.1). The dotted line indicates theoretical scaling factor $\frac{(k+1)(k+2)}{2}$ for derivatives up to order $k$. We find that both the forward-mode AD method and the impurity-tensor method closely follow the expected scaling, indicating that their bottleneck costs are essentially identical, as expected theoretically.
Figure 3 shows the total elapsed time of HOTRG computations up to the first and second derivatives at volume $V=2^{40}$ with $D=20$ to $70$, measured on a single CPU. The vertical axis shows the elapsed time of each method normalized by the execution time of the original HOTRG without derivatives. The dotted line again represents theoretical bottleneck scaling factor $\frac{(k+1)(k+2)}{2}$.
For the first-derivative case, the total runtime of the forward-mode AD method is comparable to that of the impurity-tensor method. In the two-dimensional HOTRG with the squeezer formulation, each coarse-graining step involves three SVDs, and hence the subleading costs are non-negligible compared to the bottleneck $O(D^{7})$ cost. As a result, the observed normalized runtime is smaller than the ideal bottleneck scaling. Nevertheless, the forward-mode AD and impurity-tensor methods remain quantitatively similar.
For the second-derivative case, the forward-mode AD method becomes slightly more expensive in practice, but it still provides comparable results. These deviations arise from the additional overhead of computing the squeezers and their derivatives. Overall, the forward-mode AD exhibits comparable computational scaling and, for the bottleneck contraction, incurs essentially the same computational cost as the impurity method, while achieving significantly higher accuracy.
@@
Figure 4: Relative error of the internal energy with respect to the exact result as a function of temperature. Circles show the forward-mode AD data for various bond dimensions. The impurity limit corresponds to setting the SVD derivatives to zero and using the bond-based initial tensor described in Sec. II.5.2; these data are shown by square symbols. The multi-impurity data from Ref. [60] are also shown by black symbols. The dotted line indicates the exact critical temperature.
@@ III.2 Some applications
Figure 6: $D$ dependence of the critical exponent $1/\nu$ obtained from the forward-mode AD BWTRG (circles). For comparison, the results from the scaling-collapse analysis in Ref. [60] are shown by black symbols.
@@ IV Discussions and summary
In this paper, we have proposed a simple framework for incorporating automatic differentiation into tensor renormalization group methods based on forward-mode AD. In our formulation, the memory cost for derivatives up to order $k$ scales as $k$ times that of the original TRG, independent of the graph depth while the contraction cost increases by a factor $(k+1)(k+2)/2$. The proposed forward-mode AD framework has been shown to be applicable to various TRG schemes, including HOTRG and BWTRG. In the limit where derivatives of the SVD are neglected, a theoretical correspondence to impurity methods can be established. In numerical calculations, derivatives of the partition function, including the internal energy and specific heat, have been accurately evaluated using the forward-mode AD formulation. In the impurity limit, consistency with established impurity approaches for both HOTRG and BWTRG is confirmed. Moreover, forward-mode AD achieves significantly higher accuracy than the impurity method, which neglects the parameter dependence of the isometric tensor [60]. We have also shown that, for the bottleneck tensor contraction in each coarse-graining step of the HOTRG, the computational cost for evaluating derivatives up to order $k$ increases by a factor of $(k+1)(k+2)/2$, while total cost of the HOTRG is slightly higher than that of the impurity method, due to the additional evaluations associated with differentiating the SVD or the squeezers. In Appendix C.1, we have shown that the same scaling behavior can be achieved for general tensor networks.
@@ Acknowledgements.
We would like to thank Shinichiro Akiyama, Jutho Haegeman, Atsushi Ueda, Adwait Naravane for fruitful discussions. Y.S. is supported by Graduate Program on Physics for the Universe (GP-PU), Tohoku University and JSPS KAKENHI Grant 25KJ0537. numerical calculations we performed on the supercomputer Pegasus under the Multidisciplinary Cooperative Research Program in CCS, University of Tsukuba.
@@ C.1 General case
In this section, we provide details of a contraction tree based differentiation method. While contraction tree based reverse mode automatic differentiation algorithms are described in Ref. [85], our method provides a general framework to compute arbitrary $k$-th derivatives of tensor networks by forward mode automatic differentiation.
@@
Figure 8: (a) The contraction tree for Eq. (63). (b) Modified contraction tree fr evaluating derivatives up to second order.

### v2

@@ Abstract
We propose a forward-mode automatic differentiation (AD) framework for tensor renormalization group (TRG) methods. In this approach, evaluating the derivatives of the partition function up to order $k$ increases the matrix-multiplication cost by a factor of $(k+1)(k+2)/2$ compared to computing the free energy alone, and the memory footprint is only $k+1$ times that of the original calculation. In the limit where the derivatives of the SVD are neglected, we establish a theoretical correspondence between our forward-mode AD and conventional impurity methods. Numerically, we find that the proposed AD algorithm can calculate internal energy and specific heat significantly higher accuracy than the impurity method at comparable computational cost. We also provide a practical procedure to extract critical exponents from derivatives of the renormalized tensor in TRG calculations in both two and three dimensions. In addition, we discuss how to efficiently differentiate an arbitrary tensor network.
@@ II.3 Differentiation of the HOTRG
The derivatives of E,FE,\,F can be obtained through a differentiation of the SVD formula, which we leave for appendix (B). In the practical calculation, we adopt a Lorentzian broadening $1/x\rightarrow x/(x^{2}+\eta)$ for derivatives of isometries following Ref. [59] to avoid divergence when there are degenerate singular values. In this work, we will use improved projectors refereed as squeezers [72, 73, 49] in the HOTRG calculations. In the Ising model which we use in this work, $\eta\rightarrow\infty$ limit corresponds to the vanishing limit of the derivatives of squeezers [^1] This is because initial tensor of Ising model has reflection symmetry, which yields squeezers to be isometry itself. In general squeezers, $\eta\rightarrow\infty$ doesn’t mean $\dot{E}=\dot{F}=0$ without taking derivatives of singular values to be zero.. The above procedure completes the formulation of the first-order response within the forward-mode scheme. The extension to second-order derivatives is completely analogous. By differentiating Eq. (15) once more with respect to $\beta$, we obtain the evolution equation for the second derivative $\ddot{T}^{(n+1)}$ as follows. $T^{(n)}$, $\dot{T}^{(n)}$, and $\ddot{T}^{(n)}$.
@@
Higher-order derivatives follow straightforwardly by iterating the same procedure. The update rule for $\dot{T}$ in Eq. (15) naively consists of four HOTRG-like contractions. However, we do not need to evaluate these four contractions separately. Instead, we use the fact that each contraction can be decomposed into a sequence of matrix multiplications along a contraction tree [74]. Rather than evaluating the four contractions independently, we differentiate the intermediate tensors that appear along this tree. With this procedure, the total matrix multiplication cost of computing derivatives up to first order is exactly three times that of the original HOTRG.
@@
The blue colored tensors indicate its second derivative. In this contraction, the bottleneck takes just three times as the HOTRG. In practice, it is not necessary to execute all contractions shown in the figures independently. Instead, by performing the bulk, first-order, and second-order contractions sequentially for each intermediate tensor, one can reuse the intermediate tensors that are common to Eqs. (II.1), (II.3), and (II.3), rather than recomputing them multiple times. As a result, the total cost of evaluating derivatives up to first and second order is exactly three and six times that of the original HOTRG, respectively. This scaling is unchanged for general tensor network contractions. See Appendix C.1 for details. We also note that the memory footprint is only $k+1$ times that of the original HOTRG, since we only propagate $T^{(i)}$ and its derivatives at every coarse-graining step; importantly, it does not depend on the graph depth $\log V$. Note that, to obtain the derivatives of the squeezers (isometries), we differentiate the SVD and apply the resulting rules to all contractions appearing in the original algorithm. This approach is applicable to a wide range of programming languages without relying on external AD packages, as it only requires implementing a simple update rule.
@@ III Numerical results
Next, we discuss the computational cost. We first measure the elapsed time of the bottleneck contraction part in Eq. (II.1) and its derivatives, or impurity update rule. Figure 2 shows the elapsed time of the bottleneck contraction in HOTRG for the forward-mode AD computations up to first- and second-order, and for the corresponding first- and second-order impurity tensor computations. These correspond to Eqs. (II.1), (II.3), and (II.3) for the forward-mode AD, and to Eqs. (II.1), (30), and (31) for the impurity method. For the forward-mode AD, we employ contraction tree based differentiation described in Appendix C.1. All benchmarks are performed on a single CPU using randomly generated dense tensors, and the bond dimension is varied from $D=20$ to $70$. The vertical axis shows the elapsed time normalized by that of the original HOTRG contraction in Eq. (II.1). The dotted line indicates theoretical scaling factor $\frac{(k+1)(k+2)}{2}$ for derivatives up to order $k$. We find that both the forward-mode AD method and the impurity-tensor method closely follow the expected scaling, indicating that their bottleneck costs are essentially identical, as expected theoretically.
Figure 3 shows the total elapsed time of HOTRG computations up to the first and second derivatives at volume $V=2^{40}$ with $D=20$ to $70$, measured on a single CPU. The vertical axis shows the elapsed time of each method normalized by the execution time of the original HOTRG without derivatives. The dotted line again represents theoretical bottleneck scaling factor $\frac{(k+1)(k+2)}{2}$. In the two-dimensional HOTRG with the squeezer formulation, each coarse-graining step involves three SVDs, and hence the subleading costs are non-negligible compared to the bottleneck $O(D^{7})$ cost. As a result, the observed normalized runtime is smaller than the ideal bottleneck scaling. For both first- and second-derivative case, the forward-mode AD method becomes slightly more expensive in practice, but it still provides comparable results. These deviations arise from the additional overhead of computing the squeezers and their derivatives.
Overall, the forward-mode AD exhibits comparable computational scaling and, for the bottleneck contraction, incurs essentially the same computational cost as the impurity method, while achieving significantly higher accuracy.
@@
Figure 4: Relative error of the internal energy with respect to the exact result as a function of temperature. Circles show the forward-mode AD data for various bond dimensions. The impurity limit corresponds to setting the SVD derivatives to zero and using the bond-based initial tensor described in Sec. II.5.2; these data are shown by square symbols. The multi-impurity data from Refs. [60, 78] are converted into relative errors and shown as black symbols. The dotted line indicates the exact critical temperature.
@@ III.2 Some applications
Figure 6: $D$ dependence of the critical exponent $1/\nu$ obtained from the forward-mode AD BWTRG (circles). For comparison, the results from the scaling-collapse analysis taken from Ref. [60, 78] are shown by black symbols.
@@ IV Discussions and summary
In this paper, we have proposed a simple framework for incorporating automatic differentiation into tensor renormalization group methods based on forward-mode AD. In our formulation, the memory cost for derivatives up to order $k$ scales as $k+1$ times that of the original TRG, independent of the graph depth while the contraction cost increases by a factor $(k+1)(k+2)/2$. The proposed forward-mode AD framework has been shown to be applicable to various TRG schemes, including HOTRG and BWTRG. In the limit where derivatives of the SVD are neglected, a theoretical correspondence to impurity methods can be established. In numerical calculations, derivatives of the partition function, including the internal energy and specific heat, have been accurately evaluated using the forward-mode AD formulation. In the impurity limit, consistency with established impurity approaches for both HOTRG and BWTRG is confirmed. Moreover, forward-mode AD achieves significantly higher accuracy than the impurity method, which neglects the parameter dependence of the isometric tensor [60]. We have also shown that, for the bottleneck tensor contraction in each coarse-graining step of the HOTRG, the computational cost for evaluating derivatives up to order $k$ increases by a factor of $(k+1)(k+2)/2$, while total cost of the HOTRG is slightly higher than that of the impurity method, due to the additional evaluations associated with differentiating the SVD or the squeezers. In Appendix C.1, we have shown that the same scaling behavior can be achieved for general tensor networks.
@@ Acknowledgements.
We would like to thank Shinichiro Akiyama, Jutho Haegeman, Atsushi Ueda, Adwait Naravane for fruitful discussions. Y.S. is supported by Graduate Program on Physics for the Universe (GP-PU), Tohoku University and JSPS KAKENHI Grant 25KJ0537. Numerical calculations were performed on the supercomputer Pegasus under the Multidisciplinary Cooperative Research Program in CCS, University of Tsukuba.
@@ C.1 General case
In this section, we provide details of a contraction tree based differentiation method. A contraction tree [74] is a binary tree that specifies the contraction order of a tensor networks. While contraction tree based reverse mode automatic differentiation algorithms are described in Ref. [89], our method provides a general framework to compute arbitrary $k$-th derivatives of tensor networks by forward mode automatic differentiation.
@@
Figure 8: (a) The contraction tree for Eq. (63). (b) Modified contraction tree for evaluating derivatives up to second order.